# TextInPlace: Indoor Visual Place Recognition in Repetitive Structures with Scene Text Spotting and Verification

### v1

@@ V-C Ablation Study
Figure 5: Examples of retrieval results for challenging queries on our dataset. Each column corresponds to one query case. The first three columns are from the Maze-with-Text dataset, and the last three columns are from the TextLCD dataset. The first row displays the query image, while the other two rows show the top-1 retrieved images. Incorrect retrievals are distinguished by a red border, whereas correct retrievals are highlighted with a green border. Below each image is the discriminative scene text extracted from the image.
In Table V, we conduct ablation studies to evaluate the impact of re-ranking and fine-tuning in our method. Experimental results indicate that under a repetitive structure, text-based re-ranking effectively mitigates perceptual confusion caused by similar appearances. Indoor domain fine-tuning enhances the ability of our model to learn structured features specific to indoor environments. Notably, the combination of these two components leverages their individual strengths, resulting in robust and improved performance.
@@ V-D Complexity Analysis
In Table VI, we compare the computational cost of our proposed method with the baseline method (BoQ + DG-Bridge Spotter) on our all floors dataset. In contrast to our approach, the baseline method employs BoQ [26] for image retrieval and DG-Bridge Spotter [31] for text-based re-ranking as separate processes. Our method achieves notable efficiency improvements, particularly in extraction time, which is reduced from 7.7 ms to 5.9 ms (a 23.4% reduction), and STS time, which decreases from 165.7 ms to 161.0 ms (a 2.8% reduction). These results underscore the effectiveness of our local parameter sharing network in minimizing computational latency. In the future, we can replace the current text detection and recognition modules with more lightweight alternatives to further reduce the time consumption of STS. Additionally, TextInPlace-L incurs significantly higher time consumption during the re-ranking stage compared to TextInPlace, primarily due to its reliance on remote LLM access.
@@ V-E Visualization
Fig. 5 shows detailed cases where repetitive structures cause significant perceptual confusion. When relying solely on visual descriptors, global retrieval fails in the top-1 results. However, by applying text-based verification, our method successfully finds the correct reference image. This demonstrates the effectiveness of our approach in using text-based verification to alleviate the perceptual confusion caused by repetitive structures. Furthermore, our dataset not only includes scene text but also captures the environmental structure, making it a more accurate reflection of real-world robot navigation scenarios compared to TextLCD.
@@ VI CONCLUSIONS
To address the challenge of VPR in repetitive indoor environments, TextInPlace is proposed to enable coarse-grained image retrieval using global descriptors, followed by spatial verification through discriminative scene text. Extensive experiments conducted on the newly proposed Maze-with-Text dataset, as well as the publicly available TextLCD dataset, demonstrate that the proposed TextInPlace can achieve state-of-the-art performance against extreme perceptual aliasing caused by repetitive indoor structures. Our open-source code and dataset will be available for the community’s benefit.
@@ ACKNOWLEDGMENT
The authors would like to thank Jin et al. for the contribution of TextLCD dataset and for allowing its academic use for the benefit of the community. This work was supported in part by the Shenzhen Key Laboratory of Robotics and Computer Vision (ZDSYS20220330160557001).

### v2

@@ V-C1 Selection of Fine-tuning Dataset
Figure 5: Examples of retrieval results for challenging queries. Each column corresponds to one query case. The first three columns are from the Maze-with-Text dataset, and the last three columns are from the TextLCD dataset. The first row displays the query image, while the other two rows show the top-1 retrieved images. Incorrect retrievals are distinguished by a red border, whereas correct retrievals are highlighted with a green border. Below each image is the discriminative scene text extracted from the image.
V-C2 Effectiveness of Re-ranking and Fine-tuning
As shown in Table V, We conduct ablation studies to evaluate the impact of re-ranking and fine-tuning in our method. Experimental results indicate that under a repetitive structure, text-based re-ranking effectively mitigates perceptual confusion caused by similar appearances. Indoor domain fine-tuning enhances the ability of our model to learn structured features specific to indoor environments. Notably, the combination of these two components leverages their individual strengths, resulting in robust and improved performance.
@@ V-D Complexity Analysis
The complexity of the network plays a vital role in practical applications. In Table VI, we compare the computational cost of our proposed method with the baseline method (BoQ + DG-Bridge Spotter) on our all floors dataset. In contrast to our approach, the baseline method employs BoQ [30] for image retrieval and DG-Bridge Spotter [35] for text-based re-ranking as separate processes. Our method achieves notable efficiency improvements, particularly in extraction time and STS time. These results underscore the effectiveness of our local parameter sharing network in minimizing computational latency. In the future, we can replace the text detection and recognition modules with more lightweight alternatives to further reduce the time consumption of STS. Additionally, TextInPlace-L incurs significantly higher time consumption during the re-ranking stage compared to TextInPlace, primarily due to its reliance on remote LLM access.
@@ V-E Visualization
Fig. 5 shows detailed cases where repetitive structures cause significant perceptual confusion. When relying solely on visual descriptors, global retrieval fails in the top-1 results. However, by applying text-based verification, our method successfully finds the correct reference image. This demonstrates the effectiveness of our approach in using text-based verification to alleviate the perceptual confusion caused by repetitive structures. Furthermore, our dataset not only includes scene texts but also captures the environmental structure, making it a more accurate reflection of real-world robot navigation scenarios compared to TextLCD.
@@ VI CONCLUSIONS
To address the challenge of VPR in repetitive indoor environments, TextInPlace is proposed to enable coarse-grained image retrieval using global descriptors, followed by spatial verification through discriminative scene texts. Extensive experiments conducted on the newly proposed Maze-with-Text dataset, as well as the publicly available TextLCD dataset, demonstrate that the proposed TextInPlace can achieve state-of-the-art performance against perceptual aliasing caused by repetitive indoor structures. Our code and dataset are available for the benefit of the community.
@@ ACKNOWLEDGMENT
The authors would like to thank Tongxing Jin, Thien-Minh Nguyen, Xinhang Xu, Yizhuo Yang, Shenghai Yuan, Jianping Li and Lihua Xie for the contribution of the TextLCD dataset [17] and for allowing its academic use.